Using the following chemical equation how many grams of calcium chloride are produced when 18.84 g of water are produced?

2HCl + Ca(OH)2 -> 2H2O + CaCl2

1
Expert's answer
2021-02-15T02:48:00-0500

Solution:

The balanced chemical equation:

Ca(OH)2 + 2HCl → CaCl2 + 2H2O

According to the equation above: n(CaCl2) = n(H2O)/2


Moles of H2O = Mass of H2O / Molar mass of H2O

The molar mass of H2O is 18.0153 g mol-1.

Hence,

n(H2O) = 18.84 g / 18.0153 g mol-1 = 1.04578 mol


n(CaCl2) = n(H2O) / 2 = 1.04578 mol / 2 = 0.52289 mol


Moles of CaCl2 = Mass of CaCl2 / Molar mass of CaCl2

Mass of CaCl2 = Moles of CaCl2 × Molar mass of CaCl2

The molar mass of CaCl2 is 110.98 g mol-1.

Hence,

m(CaCl2) = 0.52289 mol × 110.98 g mol-1 = 58.0303 g = 58.0 g


Answer: 58.0 grams of calcium chloride (CaCl2) are produced.